1. (x-1)(x+1/2-root3/2i)(x+1/2+root3/2i)=?

Grade:10

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:

Hello Student,
Please find answer to your question below

(x-1)(x+\frac{1}{2}-\frac{\sqrt{3}}{2i})(x+\frac{1}{2}+\frac{\sqrt{3}}{2i}) = 0
(x-1)(x^{2}+(\frac{1}{2}-\frac{\sqrt{3}}{2i}+\frac{1}{2}-\frac{\sqrt{3}}{2i})x+1) = 0
(x-1)(x^{2}+x+1) = 0
(x^{3}+x^2+x-x^2-x-1) = 0
x^{3} - 1